OpenStudy (anonymous):

log6(a^(2)+2)+log6(2)=2 help please

OpenStudy (anonymous):

Possible intermediate steps: (log(a^2+2))/(log(6))+(log(2))/(log(6)) = 2 Subtract (log(2))/(log(6)) from both sides: (log(a^2+2))/(log(6)) = 2-(log(2))/(log(6)) Divide both sides by 1/(log(6)): log(a^2+2) = (2-(log(2))/(log(6))) log(6) Cancel logarithms by taking exp of both sides: a^2+2 = 18 Subtract 2 from both sides: a^2 = 16 Take the square root of both sides: a = -4 or a = 4 Now test that these solutions are appropriate by substitution into the original equation: Check the solution a = -4: (log(a^2+2))/(log(6))+(log(2))/(log(6)) => (log(2))/(log(6))+(log(2+(-4)^2))/(log(6)) = (log(2))/(log(6))+(log(18))/(log(6)) ~~ 2. So the solution is correct. Check the solution a = 4: (log(a^2+2))/(log(6))+(log(2))/(log(6)) => (log(2))/(log(6))+(log(2+4^2))/(log(6)) = (log(2))/(log(6))+(log(18))/(log(6)) ~~ 2. So the solution is correct. Thus, the solutions are: a = -4 or a = 4